|-3x+4|+6=13 solve in positive and negative terms
Answer by algebrahouse.com(1659)

|3x + 4| + 6 = 13
|3x + 4| = 7 {subtracted 6 from both sides}
3x + 4 = 7 or 3x + 4 = -7 {set that inside the absolute value equal to ±7}
3x = 3 or 3x = -11 {subtracted 4 from both sides}
x = 1 or x = -11/3 {divided both sides by 3}

For more help from me, visit: www.algebrahouse.com